OpenStudy (anonymous):

Carrington is paid at the same rate each time he baby sits. The graph shows the rates he is paid for different times. Which equation represents the situation?

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

two points are given u have to form equation with these two points y-y1=m(x-x1) m=(y1-y2)/(x1-x2)

OpenStudy (mathmath333):

x1=4,y1=33,x2=8,y2=61

OpenStudy (anonymous):

-29 -53

OpenStudy (mathmath333):

first find m

OpenStudy (mathmath333):

m=(y1-y2)/(x1-x2)

OpenStudy (anonymous):

I am confused when every i add i get the same number

OpenStudy (mathmath333):

m=(y1-y2)/(x1-x2)=(33-61)/(4-8)=-28/-4=28/4=7

OpenStudy (mathmath333):

y-y1=m(x-x1) y-y1=7(x-x1) now put x1 and y1 in the equation

OpenStudy (anonymous):

Is it Y=7x+5

OpenStudy (anonymous):

Because that is what i have got

OpenStudy (mathmath333):

how

OpenStudy (anonymous):

Rise over run Rised over 7 then run to 5

OpenStudy (mathmath333):

correct

OpenStudy (anonymous):

yes thanks for help